Calculate number of moles in
A)7.85gm Fe (molar mass =56 gm/mol)
B)4.68 mg of Si (molar mass =28.1gm/mol)
1
Expert's answer
2015-11-19T12:22:21-0500
n=m/M

A. n = 7.85g / 56g/mol = 0.14 mol

B. n = 0.00468g / 28.1 g/mol = 0.00017